### v1

@@ Abstract
Non-orthogonal unicast multicast (NOUM) is a variant of multi-antenna multi-user communications where the users desire a shared message (multicast) in addition to their respective unique messages (unicast). The multicast rate is capped in many emerging NOUM applications, such as live-event broadcasting, location-based services and vehicular communications. Given this constraint, we experimentally show that when the user channels are highly correlated, Rate-Splitting Multiple Access (RSMA)-based NOUM can meet the multicast rate while supporting a larger unicast sum rate than conventional multi-user linear precoding (MULP)-based NOUM.
@@ I Introduction
Non-orthogonal Unicast Multicast (NOUM) is a variant of multi-antenna multi-user communications where the users desire a shared message (multicast) in addition to their respective unique messages (unicast). Such a scenario has use cases in live-event broadcasting (where the unicast messages could be social media interactions), V2X networks and AR/VR applications (where the multicast message could be location-dependent information), etc.. NOUM is a spectrally efficient way to jointly realize (physical layer) multicast and unicast in multi-antenna systems, with great potential for next-general wireless networks [1], provided adequate multicast and unicast data rates can be guaranteed.
@@
    In many NOUM applications, the multicast rate is capped (e.g. video streaming, V2X safety messages, etc. [12].). Hence, a well-motivated objective is to maximize the unicast sum rate, for instance), subject to satisfying the multicast rate.
@@ II System Model for RSMA-based NOUM
We follow the RSMA-based NOUM architecture of [10]. Consider the two-user MISO case. Let $W_{0}$ denote the multicast message, and $W_{i}~{}(i=1,2)$ the unicast message for user $i$. At the TX, each $W_{i}$ is split into common and private portions, denoted by $W_{c,i}$ and $W_{p,i}$, respectively. Let $W_{c}:=[W_{0},W_{c,1},W_{c,2}]$ denote the common message formed by concatenating the multicast message along with the common portions of each unicast message. Through an appropriately chosen Modulation and Coding Scheme (MCS) level, $W_{c}$ is transformed into a common stream, $s_{c}$. Similarly, $W_{p,1}$ and $W_{p,2}$ are separately encoded and modulated to form private streams, $s_{1}$ and $s_{2}$, respectively. All three streams are linearly precoded to form the transmit signal:
@@
where $n_{i}$ is the additive receiver noise. Upon receiving $y_{i}$, the $i$-th user first decodes $s_{c}$, while treating the interference from the $s_{1}$ and $s_{2}$ as noise. This enables the $i$-th user to recover $W_{0}$ and a part of its unicast message ($W_{c,i}$). Then, through SIC, the $i$-th user decodes $s_{i}$ to recover the rest of its unicast message ($W_{p,i}$) by treating the interference from the other private stream as noise. It is easy to see from (II) that the achievable multicast and unicast rates are a function of the precoders and MCS levels chosen for each of the three streams. We describe these next.
  • Precoder Design: Let $\hat{{\mathbf{h}}}_{i}$ denote the estimate of ${\mathbf{h}}_{i}$ at the TX. We consider weighted maximum-ratio transmission (MRT) precoding for the common stream, and zero-forcing (ZF) precoding for the private streams, i.e.,
    ${\mathbf{p}}_{c}(t)$ $=\sqrt{P_{t}(1-t)}\frac{(\hat{{\mathbf{u}}}_{1}+\hat{{\mathbf{u}}}_{2})}{\sqrt{2}}$ (3)
    ${\mathbf{p}}_{i}(t)$ $=\sqrt{P_{t}\frac{t}{2}}\hat{{\mathbf{u}}}_{i,\perp}~{}(i=1,2)$ (4)
    where $\hat{{\mathbf{u}}}_{i}$ and $\hat{{\mathbf{u}}}_{i,\perp}$ denote the unit vectors along the direction of $\hat{{\mathbf{h}}}_{i}$ and its nullspace, respectively. The parameter $t\in[0,1]$ determines the allocation of the transmit power, $P_{t}$, to the three streams.
  • MCS: An MCS level can be defined by a pair ${\mathcal{I}}=(m,r)$, where positive integer $m$ denotes the bits per modulation symbol (e.g., 2 for QPSK) and $r\in(0,1]$ denotes the code rate. Let ${\mathcal{I}}_{c}=(m_{c},r_{c})$, ${\mathcal{I}}_{1}=(m_{1},r_{1})$ and ${\mathcal{I}}_{2}=(m_{2},r_{2})$ denote the MCS levels chosen for $s_{c}$, $s_{1}$ and $s_{2}$, respectively.
@@
where the probability terms capture the rate loss due to decoding errors. In particular, ${\mathbb{P}}(\hat{W}_{c}^{(1)}=W_{c};\hat{W}_{c}^{(2)}=W_{c})$ in (5) denotes the probability that $W_{c}$ is correctly decoded at both users, while ${\mathbb{P}}(\hat{W}_{p,i}=W_{p,i})$ in (6) denotes the probability that $W_{p,i}$ is correctly decoded at the $i$-th user.
@@ III Benefits of RSMA-based NOUM
RSMA-based NOUM coincides with MULP-based NOUM when $t^{*}=1$. The RSMA-based NOUM rate region (red line) and MULP-based NOUM (blue line) coincide.
@@ IV Measurement Results
We realize a $2\times 2$ MISO scenario for RSMA-based NOUM using a pair of NI USRP-2942 software-defined radios. We use OFDM signals for the three streams, with the MCS levels chosen from Table II. For more details of our experimental setup, see [9, Section III].
We consider four cases as shown in Fig. 2, and described below:
@@
  • For several values of $t$ across Figs. 3a-d, the measured multicast and unicast throughputs are below the time-sharing line. This is due to the limited granularity of the MCS levels. As an example, consider the point corresponding to $t=0.5$ for Case 2 in Fig. 3b. One way to push this point above the time-sharing line is by increasing the multicast throughput using a higher MCS level for the common stream. However, from Fig. 4b, the next highest MCS level for the common stream ($\mathcal{I}_{C}=2$) is too aggressive, causing imperfect SIC and error propagation. More broadly, points 3 and 4 underscore the importance of the dual assumptions of perfect SIC and MCS levels of arbitrary granularity in Remark 4.

### v2

@@ Abstract
Non-orthogonal unicast multicast (NOUM) is a variant of multi-antenna multi-user communications where the users desire a shared message (multicast) in addition to their respective unique messages (unicast). The multicast rate is capped in many emerging NOUM applications, such as live-event broadcasting, location-based services and vehicular communications. Given this constraint, we experimentally show that when the user channels are highly correlated, Rate-Splitting Multiple Access (RSMA)-based NOUM can meet the multicast rate while supporting a larger unicast sum rate than multi-user linear precoding (MULP)-based NOUM and orthogonal unicast multicast.
@@ I Introduction
Non-orthogonal Unicast Multicast (NOUM) is a unique form of multi-antenna multi-user communications where users desire both multicast (shared) and unicast (unique) messages. These could arise in live-event broadcasting (where the unicast messages could be social media interactions), Vehicle-to-Everything (V2X) networks and Augmented Reality/Virtual Reality (AR/VR) applications (where the multicast message could be location-dependent information), etc. NOUM is a spectrally efficient way to jointly realize (physical layer) multicast and unicast in multi-antenna systems, with great potential for next-general wireless networks [1], provided adequate multicast and unicast data rates can be guaranteed.
@@
    In many NOUM applications, the multicast rate is capped (e.g., video streaming, V2X safety messages, etc. [12]). Hence, a well-motivated objective is to maximize the unicast sum rate, subject to satisfying the multicast rate.
@@ II System Model for RSMA-based NOUM
We follow the RSMA-based NOUM architecture of [10]. Consider the two-user multiple-input single-output (MISO) case. Let $W_{0}$ denote the multicast message, and $W_{i}~{}(i=1,2)$ the unicast message for user $i$. At the TX, each $W_{i}$ is split into common and private portions, denoted by $W_{c,i}$ and $W_{p,i}$, respectively (in general, the splitting ratio need not be 50-50 nor does it have to be the same across both users). Let $W_{c}:=[W_{0},W_{c,1},W_{c,2}]$ denote the common message formed by concatenating the multicast message along with the common portions of each unicast message. Through an appropriately chosen Modulation and Coding Scheme (MCS) level, $W_{c}$ is transformed into a common stream, $s_{c}$. Similarly, $W_{p,1}$ and $W_{p,2}$ are separately encoded and modulated to form private streams, $s_{1}$ and $s_{2}$, respectively. All three streams are linearly precoded to form the transmit signal:
@@
where $n_{i}$ is the receiver thermal noise. Upon receiving $y_{i}$, the $i$-th user first decodes $s_{c}$, while treating the interference from the $s_{1}$ and $s_{2}$ as noise. This enables the $i$-th user to recover $W_{0}$ and a part of its unicast message ($W_{c,i}$). Then, through SIC, the $i$-th user decodes $s_{i}$ to recover the rest of its unicast message ($W_{p,i}$) by treating the interference from the other private stream as noise. It is easy to see from (II) that the achievable multicast and unicast rates are a function of the precoders and MCS levels chosen for each of the three streams. We describe these next.
  • Precoder Design: Let $\hat{{\mathbf{h}}}_{i}$ denote the estimate of ${\mathbf{h}}_{i}$ at the TX. Since $s_{c}$ must be decoded by both users, ${\mathbf{p}}_{c}$ must have sufficient gain along each $\hat{{\mathbf{h}}}_{i}$. Thus, weighted maximum-ratio transmission (MRT) precoding is a reasonable good heursitic choice for ${\mathbf{p}}_{c}$, i.e.,
    ${\mathbf{p}}_{c}(t)$ $:=\sqrt{P_{t}(1-t)}\frac{(\hat{{\mathbf{u}}}_{1}+\hat{{\mathbf{u}}}_{2})}{\|\hat{{\mathbf{u}}}_{1}+\hat{{\mathbf{u}}}_{2}\|},$ (3)
    where $\hat{{\mathbf{u}}}_{i}$ denotes the unit vector along $\hat{{\mathbf{h}}}_{i}$. On the other hand, user $i$ decodes $s_{i}$ by treating the interference from $s_{j}~{}(j\neq i)$ as noise. Hence, to suppress this interference as much as possible, we consider zero-forcing (ZF) precoding for the private streams, i.e.,
    ${\mathbf{p}}_{i}(t)$ $:=\sqrt{P_{t}(t/2)}\hat{{\mathbf{u}}}_{j}^{\perp}~{}(j\neq i;i=1,2),$ (4)
    where ${\mathbf{u}}_{i}^{\perp}$ denotes the unit vector along the nullspace of $\hat{{\mathbf{h}}}_{i}$. In (3) and (4), the parameter $t\in[0,1]$ determines the allocation of the transmit power, $P_{t}$, to the three streams.
  • MCS: An MCS level can be defined by a pair ${\mathcal{I}}=(m,r)$, where positive integer $m$ denotes the bits per modulation symbol (e.g., 2 for QPSK) and $r\in(0,1]$ denotes the code rate. In general, the MCS level of each stream can be distinct. Hence, let ${\mathcal{I}}_{c}=(m_{c},r_{c})$, ${\mathcal{I}}_{1}=(m_{1},r_{1})$ and ${\mathcal{I}}_{2}=(m_{2},r_{2})$ denote the MCS levels chosen for $s_{c}$, $s_{1}$ and $s_{2}$, respectively.
@@
where the probability terms capture the rate loss due to decoding errors. In particular, ${\mathbb{P}}(\hat{W}_{c}^{(1)}=W_{c};\hat{W}_{c}^{(2)}=W_{c})$ in (5) denotes the probability that $W_{c}$ is correctly decoded at both users, while ${\mathbb{P}}(\hat{W}_{p,i}=W_{p,i})$ in (6) denotes the probability that $W_{p,i}$ is correctly decoded at the $i$-th user. These probabilities are difficult to characterize in closed-form, but can be evaluated empirically (see Section IV).
@@ III Benefits of RSMA-based NOUM
When $t^{*}=1$, RSMA-based NOUM switches to unicast only (no common stream), and RSMA-based NOUM coincides with MULP-based NOUM. The RSMA-based NOUM region (red line) and MULP-based NOUM (blue line) coincide.
@@ IV Measurement Results
We realize a $2\times 2$ MISO scenario for RSMA-based NOUM using a pair of NI USRP-2942 software-defined radios. We use OFDM signals for the three streams, with the MCS levels chosen from Table II. For details of our implementation (e.g., channel estimation, OFDM frame structure, etc.), see [9]. We consider four cases as shown in Fig. 2, and described below:
@@
  • 5.
    For several values of $t$ across Figs. 3a-d, the measured multicast and unicast throughputs are below the time-sharing line. This is due to the limited granularity of the MCS levels. As an example, consider the point corresponding to $t=0.5$ for Case 2 in Fig. 3b. One way to push this point above the time-sharing line is by increasing the multicast throughput using a higher MCS level for the common stream. However, from Fig. 4b, the next highest MCS level for the common stream ($\mathcal{I}_{C}=2$) is too aggressive, causing imperfect SIC and error propagation. More broadly, points 4 and 5 underscore the importance of the dual assumptions of perfect SIC and MCS levels of arbitrary granularity in Remark 4.
@@ V Summary
VI Acknowledgment
This work was partially funded by UKRI Impact Acceleration Account (IAA) grant EP/X52556X/1 and UKRI grants EP/X040569/1, EP/Y037197/1, EP/X04047X/1, EP/Y037243/1.